# A Review of Online Diffusion Policy RL Algorithms for Scalable Robotic Control

### v1

@@ 1 Introduction
The advent of diffusion policies [diffusionpolicy, mpd, diffuser] has marked a breakthrough in Physical AI, demonstrating superior performance over conventional policy networks [agostini2010reinforcement, sql] in complex robotic manipulation tasks. This remarkable advancement stems from their inherent capacity to model multimodal action distributions [diffusionpolicy], which provides the expressive power necessary to capture the diverse and geometrically constrained action spaces required for real-world tasks. While diffusion policies excel in modeling complex and diverse robot behaviors, such as navigation [nomad, navdp], manipulation [diffusionpolicy, hdp], and locomotion tasks [diffuseloco, ze2025generalizable], their foundation in imitation learning inherently limits their potential. Specifically, policies derived via imitation are primarily bound by the optimality of the provided demonstration data [dppo], creating a performance ceiling that prevents the agent from autonomously discovering actions superior to expert demonstrations.
To address this optimality gap, recent research has explored combining offline Reinforcement Learning (offline RL) principles with diffusion policies [diffusionql, consistencyac]. This mechanism achieves superior performance and learning efficiency by guiding the policy to prioritize high-reward actions extracted from potentially sub-optimal demonstrations. Nevertheless, offline RL methods still remain constrained by the fixed nature of static datasets, rendering them incapable of discovering novel optimal actions or generalizing to out-of-distribution (OOD) scenarios. This inability to adapt to unseen states presents a significant challenge for robotic tasks, which require robustness and strict safety guarantees when deployed in dynamic, real-world environments.
To overcome the lack of autonomous exploration in offline settings, there is a growing necessity for Online Diffusion Policy RL (Online DPRL) [dipoerror, qvpoerror, dacer, genpo], which integrates diffusion policies with online RL [sac, ppo, dsac] to learn through direct interaction with the environment. Recent advancements in high-fidelity simulation platforms [mujoco, brax, isaaclab], which leverage photorealistic rendering and high physical accuracy to reduce the domain gap to the real world significantly, have fueled the feasibility of online learning for diffusion models. Moreover, the multi-environment parallelization capabilities of these simulators mitigate the multi-step sampling inherent to diffusion models by accelerating data collection. This convergence of advanced simulation and continuous policy refinement creates a compelling case for the Online DPRL paradigm. The distinct advantages of online RL over its offline and imitation learning counterparts include:
@@
Despite these potential advantages, the integration of diffusion policies with online RL remains relatively underexplored due to fundamental structural conflicts. The primary challenge lies in the incompatibility between modern RL mechanisms, which often require backpropagation through the policy network, and the surrogate objectives of diffusion models [ddpm, sgm, fmn]. First, propagating gradients through the entire reverse diffusion chain—similar to standard policy gradient methods [policygradient]—is computationally prohibitive and suffers from gradient instability such as vanishing or exploding gradients [bengio1994learning]. Second, unlike tractable Gaussian policies [ppo, agostini2010reinforcement] commonly used in modern RL, diffusion policies lack an analytical expression for likelihood or entropy, making the calculation of tractable objectives difficult. To tackle these problems, a few nascent online DPRL approaches employ various strategies such as decoupling policy improvement [dipoerror, DDiffPG] or utilizing surrogate objective reweighting [qvpoerror, dpmdsdac].
@@
This paper specifically focuses on online model-free reinforcement learning approaches that employ diffusion-based policy representations. These online model-free methods offer greater generality across domains without requiring learned dynamics models, making them more suitable for deployment in diverse robotic scenarios. We establish the theoretical foundations of RL and diffusion policies in Sec. 2, then propose a comprehensive taxonomy in Sec. 3 that categorizes existing Online DPRL methods into four distinct groups—Action-gradient, Q-weighting, Proximity-based, and BPTT-based methods—critically examining how each approach circumvents the intractability of policy improvement. Sec. 4 presents an empirical analysis within a unified NVIDIA Isaac Lab benchmark [isaaclab], which is a GPU-accelerated robotics simulation framework, integrating both classical RL algorithms [sac, ppo] and representative Online DPRL methods [dipoerror, qvpoerror, genpo, dacer] to investigate generalization, efficiency, and robustness across diverse robotic scenarios. Finally, Sec 5 synthesizes our findings to provide a practical rationale for algorithm selection under specific operational constraints, while identifying current computational bottlenecks and outlining promising future research directions to overcome these limitations. A schematic overview of the paper’s structure and its central pillars is presented in Fig. 2.
@@
2 Background
In this section, we provide a brief introduction to fundamental reinforcement learning concepts and standard model-free online RL algorithms to establish the foundation for discussing online diffusion policy RL literature (Sec 2.1). We also present a mathematical formulation of diffusion policies in RL environments (Sec 2.2). For clarity, we unify the mathematical notation as presented in Tab. I.
2.1 Reinforcement Learning Fundamentals
@@ 2.1.1 Value Optimization in RL
Despite its solid mathematical foundations, reinforcement learning struggled with approximating Q-functions in high-dimensional state spaces. Deep Q-Networks (DQN) [dql] addressed this challenge by employing neural networks as Q-function approximators, thereby enabling effective learning in high-dimensional state spaces with discrete action spaces.
2.1.2 Policy Gradient Methods
However, in tasks requiring continuous action spaces, directly solving Eq. 2 is computationally intractable because finding the global maximum over a continuous domain at every step is prohibitively expensive. This necessitates the use of parameterized policy networks $\pi_{\theta}$ and policy gradient methods [suttonbible]. Policy gradient methods directly optimize the expected return $J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[G]$ via gradient ascent on the policy parameters $\theta$. The foundational Policy Gradient Theorem [policygradient, gae] enables gradient estimation using sampled trajectories $\tau=\{s_{0},a_{0},r_{0},s_{1},a_{1},\cdots\}$ as follows:
@@ 2.1.2 Policy Gradient Methods
2.1.3 Taxonomy of Deep Reinforcement Learning
@@
Standard RL represents the most prevalent approach adopted by many foundational methods [dql, ddpg, ppo]. To maximize the expected return $G$, this approach leverages the Bellman equation to learn the Q-function by minimizing the Temporal Difference (TD) error between the predicted action-value $Q_{\phi}(s,a)$ and the target action-value $\hat{Q}(s,a)$:
@@
Maximum Entropy RL [sac, sql] augments the standard objective by incorporating a weighted policy entropy term $\beta\log\pi(a|s)$ to encourage exploration and prevent premature convergence to suboptimal policies. The action-value function is learned by minimizing the residual of the Soft Bellman Eq. [sql]:
@@
In contrast to standard approaches, Distributional RL [dsac, c51] shifts the focus from estimating the expected return $Q(s,a)$ to modeling the complete probability distribution of returns $Z(s,a)$, which satisfies $Q(s,a)=\mathbb{E}_{Z}[Z(s,a)]$. This enables the agent to reason about risk, uncertainty, and potential extreme outcomes. The return distribution $Z_{\phi}(s,a)$ is learned by minimizing the distance between the predicted distribution and the target distribution $\hat{Z}(s,a)$:
@@
where $d_{\text{Dist}}$ is a distributional distance metric, such as the Kullback-Leibler divergence [kldivergence] or the Wasserstein distance [wasersteindistance].
2.2 Diffusion Policies in Reinforcement Learning
Standard RL implementations model policies as unimodal Gaussians or deterministic mappings [ppo, ddpg, sac], fundamentally limiting their ability to represent multimodal behavior distributions required for complex robotic tasks. While Gaussian Mixture Models [agostini2010reinforcement] and energy-based policies [sql, florence2022implicit] attempted to address this, they suffer from training instability or high inference costs. Diffusion Policies [diffusionpolicy], formulating policies as conditional denoising processes, have emerged as a powerful alternative offering high expressiveness and training stability. They have demonstrated remarkable success in visuomotor control [hdp, diffuseloco, ze2025generalizable], effectively handling high-dimensional observations and actions.
2.2.1 Action Generation via Diffusion Sampling
@@
where $k$ denotes the diffusion timestep decreasing from $K$ to 0, $\epsilon_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ represents Gaussian noise scaled by $\sigma_{k}\in\mathbb{R}$, and $\alpha_{k},\bar{\alpha}_{k}\in\mathbb{R}$ are coefficients determined by a predefined noise schedule [song2020denoising]. This iterative denoising process allows the agent to construct precise actions conditioned on the current environmental state $s$.
2.2.2 Training Diffusion Policies via Imitation Learning
@@
2.2.3 Fundamental Challenges of Integration between Diffusion Policy and Online RL
@@
    Optimization Instability: Backpropagating through a long chain of operations often leads to vanishing or exploding gradient problems [bengio1994learning], degrading the learning signal for early denoising steps and destabilizing policy optimization.
@@
3 Taxonomy of Online DPRL
@@
PPO [ppo] Standard on-policy MuJoCo [mujoco] PG [policygradient], TRPO [trpo], CEM [cem], A2C [a2c], ACER [acer]
DDPG [ddpg] Standard off-policy MuJoCo, Torcs [torcs] DPG [dpg], iLQG [ilqg]
TD3 [td3] Standard off-policy MuJoCo PPO, DDPG, TRPO, ACKTR [acktr], SAC
SAC [sac] MaxEnt off-policy MuJoCo, RLLab [rllab] PPO, DDPG, SQL [sql], TD3
*DIPO [dipoerror] Standard off-policy MuJoCo PPO, TD3, SAC
DDiffPG [DDiffPG] Standard off-policy D4RL [d4rl], Panda-Gym [pandagym] TD3, SAC, Diffusion-QL [diffusionql], Consistency-AC [consistencyac], RPG [rpg], DIPO
QSM [qsm] Standard off-policy DMC [dmcontrol] TD3, SAC
*QVPO [qvpoerror] Standard off-policy MuJoCo PPO, TD3, SAC, DIPO, QSM, HD [hd]
DPMD [dpmdsdac] Standard off-policy MuJoCo PPO, TD3, SAC, QSM, DIPO, DACER, QVPO, DPPO [dppo]
SDAC [dpmdsdac] MaxEnt off-policy MuJoCo PPO, TD3, SAC, QSM, DIPO, DACER, QVPO, DPPO
*GenPO [genpo] Standard on-policy Isaac Lab [isaaclab] PPO, DDPG, TD3, SAC, DACER, QVPO
FPO [fpo] Standard on-policy DMC PPO, DPPO
*DACER [dacer] Standard+MaxEnt off-policy MuJoCo PPO, TRPO [trpo], DDPG, TD3, SAC, DSAC [dsac]
DACERv2 [dacerv2] Standard+MaxEnt off-policy MuJoCo PPO, SAC, DSAC, DIPO, DIME, DACER, QVPO
CPQL [cpql] Standard off-policy MuJoCo, DMC PPO, MPO [mpo], DMPO [dmpo], D4PG [d4pg], TD3, SAC, DIPO, DreamerV3 [dreamerv3]
DIME [dime] MaxEnt off-policy MuJoCo, DMC, MyoSuite [myosuite] CrossQ [crossq], BRO [bro], QSM, DIPO, QVPO, DACER
DSAC-D [dsacd] Distributional off-policy MuJoCo, Vehicle Meeting [dsacd] PPO, TRPO, DDPG, TD3, SAC, DSAC-T [dsact], DACER
@@
Existing online model-free DPRL methods tackle the challenges outlined in Sec. 2.2.3 through diverse algorithmic strategies. We classify these approaches into four distinct families based on their core mechanism for policy improvement: (1) Action-Gradient-Based Methods (Sec. 3.1), (2) Q-Weighting Methods (Sec. 3.2), (3) Proximity-Based Methods (Sec. 3.3), and (4) Backpropagation Through Time (BPTT) (Sec. 3.4). Each family employs fundamentally different techniques to circumvent the computational and optimization difficulties inherent in training diffusion policies with online RL. In the following sections, we provide a comprehensive review organized by this taxonomy, with a particular focus on the mathematical formulations of policy improvement. Furthermore, we summarize these concepts in the following sections: Sec. 3.5 categorizes the diverse Diffusion-based RL paradigms by analyzing how diffusion models are integrated into online policy optimization. Sec. 3.6 then identifies the cross-cutting challenges of online DPRL and discusses key optimization strategies to improve computational efficiency and training stability. A consolidated summary of online model-free DPRL approaches is presented in Fig. 3 and Tab. II.
3.1 Action-Gradient-Based Methods
To apply diffusion policy to model-free online RL tasks, several pioneering approaches [dipoerror, qsm] leverage the action gradient, which provides a novel approach for policy improvement as follows:
@@ 3.1 Action-Gradient-Based Methods
3.1.1 Model-free Online RL with DIffusion POlicy (DIPO)
DIPO [dipoerror] is the pioneering work in online Diffusion Policy RL, introducing the action gradient method for policy optimization. By first improving actions through action gradients and then using these improved actions as targets for diffusion model training, DIPO avoids the computational burden of backpropagating through the entire diffusion chain.

### v2

@@ I Introduction
The advent of diffusion policies [19, 15, 44] has marked a breakthrough in Physical AI, demonstrating superior performance over conventional policy networks [4, 36] in complex robotic manipulation tasks. This remarkable advancement stems from their inherent capacity to model multimodal action distributions [19], which provides the expressive power necessary to capture the diverse and geometrically constrained action spaces required for real-world tasks. While diffusion policies excel in modeling complex and diverse robot behaviors, such as navigation [83, 14], manipulation [19, 58], and locomotion tasks [41, 106], their foundation in imitation learning inherently limits their potential. Specifically, policies derived via imitation are primarily bound by the optimality of the provided demonstration data [73], creating a performance ceiling that prevents the agent from autonomously discovering actions superior to expert demonstrations.
To address this optimality gap, recent research has explored combining offline Reinforcement Learning (offline RL) principles with diffusion policies [97, 23]. This mechanism achieves superior performance and learning efficiency by guiding the policy to prioritize high-reward actions extracted from potentially sub-optimal demonstrations. Nevertheless, offline RL methods still remain constrained by the fixed nature of static datasets, rendering them incapable of discovering novel optimal actions or generalizing to out-of-distribution (OOD) scenarios. This inability to adapt to unseen states presents a significant challenge for robotic tasks, which require robustness and strict safety guarantees when deployed in dynamic, real-world environments.
To overcome the lack of autonomous exploration in offline settings, there is a growing necessity for Online Diffusion Policy RL (Online DPRL) [103, 21, 95, 22], which integrates diffusion policies with online RL [37, 77, 59] to learn through direct interaction with the environment. Recent advancements in high-fidelity simulation platforms [91, 31, 62], which leverage photorealistic rendering and high physical accuracy to reduce the domain gap to the real world significantly, have fueled the feasibility of online learning for diffusion models. Moreover, the multi-environment parallelization capabilities of these simulators mitigate the multi-step sampling inherent to diffusion models by accelerating data collection. This convergence of advanced simulation and continuous policy refinement creates a compelling case for the Online DPRL paradigm. The distinct advantages of online RL over its offline and imitation learning counterparts include:
@@
Despite these potential advantages, the integration of diffusion policies with online RL remains relatively underexplored due to fundamental structural conflicts. The primary challenge lies in the incompatibility between modern RL mechanisms, which often require backpropagation through the policy network, and the surrogate objectives of diffusion models [40, 82, 51]. First, propagating gradients through the entire reverse diffusion chain—similar to standard policy gradient methods [85]—is computationally prohibitive and suffers from gradient instability such as vanishing or exploding gradients [9]. Second, unlike tractable Gaussian policies [77, 4] commonly used in modern RL, diffusion policies lack an analytical expression for likelihood or entropy, making the calculation of tractable objectives difficult. To tackle these problems, a few nascent online DPRL approaches employ various strategies such as decoupling policy improvement [103, 49] or utilizing surrogate objective reweighting [21, 57].
@@
This paper specifically focuses on online model-free reinforcement learning approaches that employ diffusion-based policy representations. These online model-free methods offer greater generality across domains without requiring learned dynamics models, making them more suitable for deployment in diverse robotic scenarios. We establish the theoretical foundations of RL and diffusion policies in Sec. II, then propose a comprehensive taxonomy in Sec. III that categorizes existing Online DPRL methods into four distinct groups—Action-gradient, Q-weighting, Proximity-based, and BPTT-based methods—critically examining how each approach circumvents the intractability of policy improvement. Sec. IV presents an empirical analysis within a unified NVIDIA Isaac Lab benchmark [62], which is a GPU-accelerated robotics simulation framework, integrating both classical RL algorithms [37, 77] and representative Online DPRL methods [103, 21, 22, 95] to investigate generalization, efficiency, and robustness across diverse robotic scenarios. Finally, Sec V synthesizes our findings to provide a practical rationale for algorithm selection under specific operational constraints, while identifying current computational bottlenecks and outlining promising future research directions to overcome these limitations. A schematic overview of the paper’s structure and its central pillars is presented in Fig. 2.
@@
II Background
In this section, we provide a brief introduction to fundamental reinforcement learning concepts and standard model-free online RL algorithms to establish the foundation for discussing online diffusion policy RL literature (Sec II-A). We also present a mathematical formulation of diffusion policies in RL environments (Sec II-B). For clarity, we unify the mathematical notation as presented in Tab. I.
II-A Reinforcement Learning Fundamentals
@@ II-A1 Value Optimization in RL
Despite its solid mathematical foundations, reinforcement learning struggled with approximating Q-functions in high-dimensional state spaces. Deep Q-Networks (DQN) [64] addressed this challenge by employing neural networks as Q-function approximators, thereby enabling effective learning in high-dimensional state spaces with discrete action spaces.
II-A2 Policy Gradient Methods
However, in tasks requiring continuous action spaces, directly solving Eq. 2 is computationally intractable because finding the global maximum over a continuous domain at every step is prohibitively expensive. This necessitates the use of parameterized policy networks $\pi_{\theta}$ and policy gradient methods [84]. Policy gradient methods directly optimize the expected return $J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[G]$ via gradient ascent on the policy parameters $\theta$. The foundational Policy Gradient Theorem [85, 76] enables gradient estimation using sampled trajectories $\tau=\{s_{0},a_{0},r_{0},s_{1},a_{1},\cdots\}$ as follows:
@@ II-A2 Policy Gradient Methods
II-A3 Taxonomy of Deep Reinforcement Learning
@@
Standard RL represents the most prevalent approach adopted by many foundational methods [64, 50, 77]. To maximize the expected return $G$, this approach leverages the Bellman equation to learn the Q-function by minimizing the Temporal Difference (TD) error between the predicted action-value $Q_{\phi}(s,a)$ and the target action-value $\hat{Q}(s,a)$:
@@
Maximum Entropy RL [37, 36] augments the standard objective by incorporating a weighted policy entropy term $\beta\log\pi(a|s)$ to encourage exploration and prevent premature convergence to suboptimal policies. The action-value function is learned by minimizing the residual of the Soft Bellman Eq. [36]:
@@
In contrast to standard approaches, Distributional RL [59, 20] shifts the focus from estimating the expected return $Q(s,a)$ to modeling the complete probability distribution of returns $Z(s,a)$, which satisfies $Q(s,a)=\mathbb{E}_{Z}[Z(s,a)]$. This enables the agent to reason about risk, uncertainty, and potential extreme outcomes. The return distribution $Z_{\phi}(s,a)$ is learned by minimizing the distance between the predicted distribution and the target distribution $\hat{Z}(s,a)$:
@@
where $d_{\text{Dist}}$ is a distributional distance metric, such as the Kullback-Leibler divergence [47] or the Wasserstein distance [93].
II-B Diffusion Policies in Reinforcement Learning
Standard RL implementations model policies as unimodal Gaussians or deterministic mappings [77, 50, 37], fundamentally limiting their ability to represent multimodal behavior distributions required for complex robotic tasks. While Gaussian Mixture Models [4] and energy-based policies [36, 30] attempted to address this, they suffer from training instability or high inference costs. Diffusion Policies [19], formulating policies as conditional denoising processes, have emerged as a powerful alternative offering high expressiveness and training stability. They have demonstrated remarkable success in visuomotor control [58, 41, 106], effectively handling high-dimensional observations and actions.
II-B1 Action Generation via Diffusion Sampling
@@
where $k$ denotes the diffusion timestep decreasing from $K$ to 0, $\epsilon_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ represents Gaussian noise scaled by $\sigma_{k}\in\mathbb{R}$, and $\alpha_{k},\bar{\alpha}_{k}\in\mathbb{R}$ are coefficients determined by a predefined noise schedule [80]. This iterative denoising process allows the agent to construct precise actions conditioned on the current environmental state $s$.
II-B2 Training Diffusion Policies via Imitation Learning
@@
II-B3 Fundamental Challenges of Integration between Diffusion Policy and Online RL
@@
    Optimization Instability: Backpropagating through a long chain of operations often leads to vanishing or exploding gradient problems [9], degrading the learning signal for early denoising steps and destabilizing policy optimization.
@@
III Taxonomy of Online DPRL
@@
PPO [77] Standard on-policy MuJoCo [91] PG [85], TRPO [75], CEM [87], A2C [63], ACER [98]
DDPG [50] Standard off-policy MuJoCo, Torcs [102] DPG [78], iLQG [89]
TD3 [33] Standard off-policy MuJoCo PPO, DDPG, TRPO, ACKTR [101], SAC
SAC [37] MaxEnt off-policy MuJoCo, RLLab [28] PPO, DDPG, SQL [36], TD3
*DIPO [103] Standard off-policy MuJoCo PPO, TD3, SAC
DDiffPG [49] Standard off-policy D4RL [32], Panda-Gym [34] TD3, SAC, Diffusion-QL [97], Consistency-AC [23], RPG [42], DIPO
QSM [71] Standard off-policy DMC [88] TD3, SAC
*QVPO [21] Standard off-policy MuJoCo PPO, TD3, SAC, DIPO, QSM, HD [17]
DPMD [57] Standard off-policy MuJoCo PPO, TD3, SAC, QSM, DIPO, DACER, QVPO, DPPO [73]
SDAC [57] MaxEnt off-policy MuJoCo PPO, TD3, SAC, QSM, DIPO, DACER, QVPO, DPPO
*GenPO [22] Standard on-policy Isaac Lab [62] PPO, DDPG, TD3, SAC, DACER, QVPO
FPO [61] Standard on-policy DMC PPO, DPPO
*DACER [95] Standard+MaxEnt off-policy MuJoCo PPO, TRPO [75], DDPG, TD3, SAC, DSAC [59]
DACERv2 [96] Standard+MaxEnt off-policy MuJoCo PPO, SAC, DSAC, DIPO, DIME, DACER, QVPO
CPQL [18] Standard off-policy MuJoCo, DMC PPO, MPO [2], DMPO [1], D4PG [8], TD3, SAC, DIPO, DreamerV3 [38]
DIME [16] MaxEnt off-policy MuJoCo, DMC, MyoSuite [13] CrossQ [11], BRO [68], QSM, DIPO, QVPO, DACER
DSAC-D [54] Distributional off-policy MuJoCo, Vehicle Meeting [54] PPO, TRPO, DDPG, TD3, SAC, DSAC-T [27], DACER
@@
Existing online model-free DPRL methods tackle the challenges outlined in Sec. II-B3 through diverse algorithmic strategies. We classify these approaches into four distinct families based on their core mechanism for policy improvement: (1) Action-Gradient-Based Methods (Sec. III-A), (2) Q-Weighting Methods (Sec. III-B), (3) Proximity-Based Methods (Sec. III-C), and (4) Backpropagation Through Time (BPTT) (Sec. III-D). Each family employs fundamentally different techniques to circumvent the computational and optimization difficulties inherent in training diffusion policies with online RL. In the following sections, we provide a comprehensive review organized by this taxonomy, with a particular focus on the mathematical formulations of policy improvement. Furthermore, we summarize these concepts in the following sections: Sec. III-E categorizes the diverse Diffusion-based RL paradigms by analyzing how diffusion models are integrated into online policy optimization. Sec. III-F then identifies the cross-cutting challenges of online DPRL and discusses key optimization strategies to improve computational efficiency and training stability. A consolidated summary of online model-free DPRL approaches is presented in Fig. 3 and Tab. II.
III-A Action-Gradient-Based Methods
To apply diffusion policy to model-free online RL tasks, several pioneering approaches [103, 71] leverage the action gradient, which provides a novel approach for policy improvement as follows:
@@ III-A Action-Gradient-Based Methods
III-A1 Model-free Online RL with DIffusion POlicy (DIPO)
DIPO [103] is the pioneering work in online Diffusion Policy RL, introducing the action gradient method for policy optimization. By first improving actions through action gradients and then using these improved actions as targets for diffusion model training, DIPO avoids the computational burden of backpropagating through the entire diffusion chain.